huff:

The three sides of a triangle are x - 1, x, and x + 1 meters. Solve for x taking the perimeter of the triangle as 18 meters.

huff:

please help!

carmelle:

You can set up this problem as an equation. (x - 1) + (x) + (x + 1) = 18 First we combine like terms. (x + x + x) + (1 - 1) = 18 3x = 18 Then slove for x by dividing both sides by 3. \[\frac{ 3x }{ 3 }=\frac{ 18 }{ 3 }\] What is 18 รท 3?